OpenStudy (anonymous):

Medal>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>If the sin 60° = square root of three over two, then which statement is true? cos 30° = square root of three over two, because the cosine and sine are complements cos 120° = square root of three over two, because the cosine and sine are supplements cos 30° = 0, because the cosine and sine are complements cos 1200° = 0, because the cosine and sine are supplements

OpenStudy (anonymous):

@Megm4301

OpenStudy (anonymous):

@NinjaDevo

OpenStudy (anonymous):

im not good at trig

OpenStudy (anonymous):

Me neither...I'm not even taking trig yet... @Kainui @Whitemonsterbunny17 @iPwnBunnies @phi Can any of you solve this?

OpenStudy (anonymous):

thanks ninja

OpenStudy (anonymous):

@Megm4301

OpenStudy (anonymous):

do you know this?

OpenStudy (anonymous):

Which pair of angles has congruent values for the sin x° and the cos y°? 10°; 170° 10°; 80° 10°; 20° 10°; 10°

OpenStudy (anonymous):

thats the other question i have

OpenStudy (anonymous):

@xiphoidx

OpenStudy (anonymous):

@precal

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@D3xt3R

OpenStudy (anonymous):

i need help with both of these questions @iGreen

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@D3xt3R

OpenStudy (anonymous):

I'm pretty sure for the the first question it would be the first option @ArmyBoy97 or at least i want to say it is.

OpenStudy (anonymous):

\[cos(30^o)=sin(60^o)=\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

The correct answer is: cos 30° = square root of three over two, because the cosine and sine are complements

OpenStudy (anonymous):

ok and the second one?

OpenStudy (anonymous):

\[Sin(10^o)=Cos(80^o)\]Always \[Sin(x)=Cos(90-x)\]

OpenStudy (anonymous):

ok i understand d3xt3r

OpenStudy (anonymous):

what about the second one

OpenStudy (anonymous):

?

OpenStudy (anonymous):

The answer is 10°; 80°

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Nop